OpenStudy (anonymous):

x to the 3rd power = 9 : rational or irrational x to the 3rd power = -343 :rational or irrational

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[x^3=9\] right?

OpenStudy (anonymous):

yep!

OpenStudy (misty1212):

then \(x=\sqrt[3]{9}\) which is not a rational number

OpenStudy (misty1212):

on the other hand if \(x^3=-343\) then \(x=\sqrt[3]{-343}\) which you can compute here http://www.wolframalpha.com/input/?i=cube+root%28-343%29